OpenStudy (anonymous):

The Indianapolis 500 is one of the most exciting events in sports. Each spring, 33 drivers compete in the 500-mile race, sometimes hitting speeds of more than 220 miles an hour. The general formula for computing distance is rate (or speed) multiplied by time, given a steady average speed. The formula may be written d = rt. Answer today's project questions on speed, distance, and time. (theres more to the problem read further)

OpenStudy (anonymous):

An auto race consists of 8 laps. A driver completes the first 3 laps at an average speed of 185 miles/hour and the remaining laps at an average speed of 200 miles/hour. If d represents the length of one lap, which expression represents the time in hours, t, that it takes the driver to complete the race?

OpenStudy (anonymous):

i got 3d/185 + 5d/200 = t

OpenStudy (anonymous):

The driver's average speed to complete all 8 laps is ______ miles per hour. Round to the nearest tenth of a mile per hour. (Enter only the number.)

OpenStudy (anonymous):

@myininaya

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

delete useless information ... The formula may be written d = rt. Answer today's project questions on speed, distance, and time. An auto race consists of 8 laps. A driver completes the first 3 laps at an average speed of 185 miles/hour and the remaining laps at an average speed of 200 miles/hour. If d represents the length of one lap, which expression represents the time in hours, t, that it takes the driver to complete the race? 3d miles * 1 hr/185 miles + 5d miles * 1 hr/200 miles

OpenStudy (amistre64):

t = 61d/1480 hours since d = rt, then d/t = r

OpenStudy (amistre64):

d/t = 1480/61 right?